OpenStudy (anonymous):

Simplify: 4x^2 y^5 / 8x^2 y

OpenStudy (lalaly):

\[\frac{4}{8} x^{2-2} y^{5-1}\]

OpenStudy (lalaly):

\[=\frac{1}{2} x^0 y^4 = \frac{1}{2}y^4\]

OpenStudy (anonymous):

0.5y^4